A grandfather clock on earth has a pendulum with a period of 1.0. Calculate the length of the pendulum


1
Expert's answer
2022-03-29T12:29:10-0400

Answer

the length of the pendulum

"L=\frac{T^2g}{4\pi}\\=\frac{1*2*9.8}{4*3.14*3.14}\\=0.50m"